OpenStudy (anonymous):

Solve for x: |2x + 6| - 4 = 20

Nnesha (nnesha):

move -4 to other side first

OpenStudy (anonymous):

oh sorry i forgot these

OpenStudy (anonymous):

1) x = 9 and x = 11 2) x = -9 and x = 15 3) x = 9 and x = -15 4) No solution

OpenStudy (texaschic101):

| 2x + 6 | - 4 = 20 --- add 4 to both sides | 2x + 6 | = 20 + 4 | 2x + 6 | = 24 2x + 6 = 24 2x = 24 - 6 2x = 18 x = 9 -(2x + 6) = 24 -2x - 6 = 24 -2x = 24 + 6 -2x = 30 x = -15 x = 9 and x = -15

OpenStudy (anonymous):

oh i keep forgetting to change the sign omg! thanks for the help

OpenStudy (texaschic101):

no problem....you got to remember about the sign